OpenStudy (anonymous):

After spending 5/9 of his salary, a man has $160 left. What amount did he spend?

Parth (parthkohli):

\[\LARGE {5 \over 9} * x = 160\] Can solve it easily now. x = 160 * 9/5 x = 38 * 9 x = 342 HIS SALARY IS $342

OpenStudy (anonymous):

160 * 9/5 is 288

OpenStudy (anonymous):

288 is his salary?

OpenStudy (anonymous):

Then what amount did he spent?

Parth (parthkohli):

Oops, yah

Parth (parthkohli):

Subtract 288 - 160. You get 128

Parth (parthkohli):

My first step went wrong.

Parth (parthkohli):

128 is the answer, right

OpenStudy (anonymous):

The answer given is 200. :S

Parth (parthkohli):

EH? lemme see it again.

Parth (parthkohli):

OH GOT IT

Parth (parthkohli):

5x / 9 = 160 is the equation

Parth (parthkohli):

5x = 160*9 5x = 1440 x = 288 It's the same

Parth (parthkohli):

So, y - 288 = 160 y = 160 + 288

Parth (parthkohli):

His salary is 448 :/

OpenStudy (anonymous):

its x-(5x/9)=160 4x/9=160 4x=160*9 x=160*9/4 x=360 so salary = 360 money spent= 360-160 (reason u r wrong is u tkng 5x/9 as money left bt its money spent)

OpenStudy (anonymous):

thank you so much.